If 1, 1/2, 0 are direction ratios of a line, then the line makes an angle of 90° with the x-axis, an angle of 60° with the y-axis, and is parallel to the z-axis. State whether true or false, also give reason for your answer.


1
Expert's answer
2021-05-04T13:14:35-0400

We have given the direction ratios "1, \dfrac{1}{2}, 0"


Also we have given the angles "90 ^\circ, 60^ \circ, 0^ \circ"


We know,

"cos\alpha = \dfrac{a}{\sqrt{a^2+b^2+c^2}}"


"cos \alpha = \dfrac{1}{\sqrt{\dfrac{5}{4}}}" ="\dfrac{2}{\sqrt5}"


"cos\beta = \dfrac{b}{\sqrt{a^2+b^2+c^2}}"


"cos \beta = \dfrac{\dfrac{1}{2}}{\sqrt{\dfrac{5}{4}}}" = "\dfrac{1}{\sqrt5}"


"cos\gamma = \dfrac{c}{\sqrt{a^2+b^2+c^2}}" = 0


Hence, the calculated values of direction cosine does not match with the given values.

Therefore, it is False .